# G3: An Effective and Adaptive Framework for Worldwide Geolocalization Using Large Multi-Modality Models

### v1

@@ Abstract
Worldwide geolocalization aims to locate the precise location at the coordinate level of photos taken anywhere on the Earth. It is very challenging due to 1) the difficulty of capturing subtle location-aware visual semantics, and 2) the heterogeneous geographical distribution of image data. As a result, existing studies have clear limitations when scaled to a worldwide context. They may easily confuse distant images with similar visual contents, or cannot adapt to various locations worldwide with different amounts of relevant data. To resolve these limitations, we propose G3, a novel framework based on Retrieval-Augmented Generation (RAG). In particular, G3 consists of three steps, i.e., Geo-alignment, Geo-diversification, and Geo-verification to optimize both retrieval and generation phases of worldwide geolocalization. During Geo-alignment, our solution jointly learns expressive multi-modal representations for images, GPS and textual descriptions, which allows us to capture location-aware semantics for retrieving nearby images for a given query. During Geo-diversification, we leverage a prompt ensembling method that is robust to inconsistent retrieval performance for different image queries. Finally, we combine both retrieved and generated GPS candidates in Geo-verification for location prediction. Experiments on two well-established datasets IM2GPS3k and YFCC4k verify the superiority of G3 compared to other state-of-the-art methods. Our code is available online  [^1] The link will be released after publication. for reproduction.
@@ 2 Related Work
Image Geolocalization. Previous work in image geolocalization can be divided into three main categories: classification-based methods, retrieval-based methods, and generation-based methods. (1) Classification-based methods seo2018cplanet ; vo2017revisiting ; muller2018geolocation ; weyand2016planet ; pramanick2022world ; clark2023we divide the entire earth into multiple grid cells and assign the center coordinates as predicted values. Models are then trained to classify the input image into the correct cell. However, if the actual location of the image is far from the center of the predicted cell, there can still be significant errors, even if the cell prediction is correct. (2) Retrieval-based methods treat the image geolocalization task as a retrieval problem, typically maintaining a database of images workman2015wide ; liu2019lending ; zhu2021vigor ; yang2021cross ; zhu2022transgeo ; tian2017cross ; shi2020looking ; zhu2023r2former or a gallery of GPS coordinates cepeda2023geoclip . They take the most similar images and GPS coordinates to the query image as the predicted values. However, maintaining a global-level image database or GPS gallery is infeasible. (3) Generation-based methods employ large multi-modality models to generate the predicted coordinates for images. Zhou et al. zhou2024img2loc introduced retrieval-augmented generation into the geolocalization task and took the retrieved similar images’ coordinates as references to help generate predictions. However, they can not accurately extract visual semantics to indicate an image’s location because they simply use visual similarity to retrieve references and suffer inaccurate prediction when facing heterogeneous query images.
Large Multi-modality Models. Inspired by the success of large models in single domains like computer vision wang2024visionllm and natural language processing zhao2023survey , there has been increasing attention on large multi-modality models. CLIP radford2021learning aligns image and text representations through contrastive learning and achieves remarkable model generalization with simple optimization objectives. The Large Language-and-Vision Assistant (LLAVA) liu2024visual effectively combines CLIP’s visual encoder with the powerful language model Vicuna, enhancing the model’s general understanding of both visual and textual information through two-stage instruction tuning. GPT4V achiam2023gpt is a large multimodal model released by OpenAI in 2023, allowing users to input text and images to obtain answers.
@@ 3.1 Database Construction
where $\mathcal{L}_{a,b}$ denotes the loss function of modality $a$ to modality $b$, $\mathbf{e}$ is the encoded representations, and $t$ is the temperature. G3 needs to align image representations with both textual descriptions and GPS data, so the final optimization objective is shown below: $\mathcal{L}=(\mathcal{L}_{\text{image},\text{text}}+\mathcal{L}_{\text{image},\text{gps}}+\mathcal{L}_{\text{text},\text{image}}+\mathcal{L}_{\text{gps},\text{image}})/4$.
@@ 3.2 Location Prediction
Geo-verification. Given the coordinate candidates set $\{c_{1},c_{2},\cdots,c_{m}\}$, selecting the best guess is essential and challenging. We reinvent the well-trained Image-to-GPS model in Geo-alignment to achieve this target, as shown in Figure 2(D). The similarity between image representations $\mathbf{e}_{\text{gps}}^{\text{image}}$ and GPS representations $\mathbf{e}^{gps}$ are derived by $\text{sim}=\mathbf{e}_{\text{gps}}^{\text{image}}(\mathbf{e}^{\text{image}})^{\text{T}}$, and the coordinate $c_{j}$ with the highest similarity is selected as the final prediction by $j=\text{argmax}(\text{sim}_{j}),\ j\in\{1,2,\cdots,m\}$.
@@ 4 MP16-Pro Dataset
To facilitate subsequent research, we propose the MP16-Pro dataset by adding textual geographical descriptions to each sample from the MediaEval Placing Tasks 2016 (MP-16) dataset larson2017benchmarking . Specifically, we utilize the open-source geocoding tool Nominatim to obtain multi-level geographical textual descriptions for each sample’s GPS location (total 4.72 million locations). There are eight geographical unit levels: neighborhood, city, county, state, region, country, country code, and continent. Some examples are given in Appendix A.1 for reference. Geographical text descriptions provide additional information for geolocalization tasks and enable models to transcend the original paradigm solely supporting image and GPS alignment, facilitating more diverse modeling approaches. MP16-Pro has been released on our repository. Please refer to Appendix A.4 to check the dataset.
@@ 5 Experiments
Implementation details: We use faiss douze2024faiss to deploy the image database. The vision and text encoders are pretrained ViT-L/14 and a masked self-attention transformer from CLIP radford2021learning . The dimensions for two trainable layers of $f_{\text{text}}$, $f_{\text{gps}}$, $f$ are 768 and 768. The input dimension of GPS encoder is 512, and the dimensions for four hidden layers of $f_{k}$ in Equation 3.1 are 1024, the output dimension is 512. For the Earth radius, we set it as 6378137.0. For RFF, we use three hierarchies with $\sigma_{min}$ as $2^{0}$ and $\sigma_{max}$ as $2^{8}$. GPT4V  [^3] https://openai.com/ is selected as the LMMs in this paper. Its temperature is set to 1.2. The number of RAG prompts $K$ is set to 4, and the number of candidates for each RAG $N$ is set to 5 for IM2GPS3K and 1 for YFCC4K. The number of similar image coordinates taken into account in candidates is 0 for IM2GPS3K and 1 for YFCC4K. G3 is trained using AdamW optimizer with learning rate 3e-5 and weight decay 1e-6. A step linear scheduler is employed with gamma 0.87, and the training epoch is set to 10. Training batch size is set to 256 and temperature $t$ in Equation 2 is initialized as 3.99. All experiments are conducted with Pytorch and one NVIDIA H800 GPU. Please refer to Appendix A.2 for more details on the training environment, training time, and API cost. We also mention the reproduction details in Appendix A.4 and the limitations of G3 in Appendix A.5.
Baselines: To evaluate G3 in geolocalization, we follow previous work cepeda2023geoclip ; zhou2024img2loc and select the following baselines for comparison: [L]kNN,$\sigma$=4 vo2017revisiting , PlaNet weyand2016planet , CPlaNet seo2018cplanet , ISNs muller2018geolocation , Translocator pramanick2022world , GeoDecoder clark2023we , GeoCLIP cepeda2023geoclip , Img2Loc zhou2024img2loc . The detailed descriptions of baselines are in Appendix A.6. Due to the lack of available implementations for Img2Loc, we reproduce it based on its paper and release it in our repository for future research.
@@ 5.1 Comparison with State-of-the-art Methods
Table 1: Overall experimental results on IM2GPS3K and YFCC4K. “ *” indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline.
@@
[L]kNN, sigma=4 7.2 19.4 26.9 38.9 55.9 2.3 5.7 11 23.5 42
PlaNet 8.5 24.8 34.3 48.4 64.6 5.6 14.3 22.2 36.4 55.8
CPlaNet 10.2 26.5 34.6 48.6 64.6 7.9 14.8 21.9 36.4 55.5
ISNs 10.5 28 36.6 49.7 66 6.5 16.2 23.8 37.4 55
Translocator 11.8 31.1 46.7 58.9 80.1 8.4 18.6 27 41.1 60.4
GeoDecoder 12.8 33.5 45.9 61 76.1 10.3 24.4 33.9 50 68.7
GeoCLIP 14.11 34.47 50.65 69.67 83.82 9.59 19.31 32.63 55 74.69
Img2Loc 15.34 39.83 53.59 69.7 82.78 19.78 30.71 41.4 58.11 74.07
Ours 16.65* 40.94* 55.56* 71.24* 84.68* 23.99* 35.89* 46.98* 64.26* 78.15*
To verify the effectiveness of G3, we conduct comparative experiments on IM2GPS3K and YFCC4K with other state-of-the-art methods. The results are shown in Table 1. (1) G3 is superior to all the other baselines on all metrics. In addition, compared to the second best methods, G3 achieves 8.5%, 2.8%, 3.7%, 2.2%, 1.0% improvements on IM2GPS3K and 21.3%, 16.9%, 13.5%, 10.6%, 4.6% improvements on YFCC4K in the 1km, 25km, 200km, 750km, 2500km thresholds. (2) G3, Img2Loc, and GeoCLIP achieve leading results, which can be attributed to the other methods taking the worldwide geolocalization task as a classification problem, introducing inevitable systemic biases. (3) G3 demonstrates significant improvements over GeoCLIP because GeoCLIP is constrained by the settings of the GPS gallery, which can not cover the entire globe. (4) Compared to Img2Loc, G3, through Geo-alignment that aligns images with discrete and continuous geographical features, achieves more precise retrieval of reference coordinates for subsequent RAG processes. Additionally, Geo-diversification and Geo-verification effectively expand the candidate pool and filter out the confident prediction results, further enhancing geolocalization performance. Overall, G3 achieves the best performance on all datasets across all metrics, which verifies the effectiveness of G3.
@@ 6 Conclusion
In this paper, we propose a novel worldwide geolocalization framework named G3. First, we introduce Geo-alignment to capture location-aware semantics in images by aligning images with textual geographical descriptions and GPS information. Second, Geo-diversification is proposed to improve the robustness of prediction generation via a prompt ensemble technique. Finally, Geo-verification selects the final coordinate prediction using the learned multi-modality representations. G3 is evaluated on two well-established datasets, IM2GPS3K and YFCC4K, and achieves state-of-the-art performance.
@@ A.1 MP16-Pro Dataset Samples
Due to size limitations, the metadata for images can not be uploaded to the anonymous repository or supplementary materials. We will release the materials as soon as possible. Additionally, downloading images based on each image URL is a time-consuming and labor-intensive process. Therefore, we have packaged the image data and will make it available for download at an appropriate time. However, this has not been achieved due to anonymity restrictions and anonymous repository storage limitations.

### v2

@@ Abstract
Worldwide geolocalization aims to locate the precise location at the coordinate level of photos taken anywhere on the Earth. It is very challenging due to 1) the difficulty of capturing subtle location-aware visual semantics, and 2) the heterogeneous geographical distribution of image data. As a result, existing studies have clear limitations when scaled to a worldwide context. They may easily confuse distant images with similar visual contents, or cannot adapt to various locations worldwide with different amounts of relevant data. To resolve these limitations, we propose G3, a novel framework based on Retrieval-Augmented Generation (RAG). In particular, G3 consists of three steps, i.e., Geo-alignment, Geo-diversification, and Geo-verification to optimize both retrieval and generation phases of worldwide geolocalization. During Geo-alignment, our solution jointly learns expressive multi-modal representations for images, GPS and textual descriptions, which allows us to capture location-aware semantics for retrieving nearby images for a given query. During Geo-diversification, we leverage a prompt ensembling method that is robust to inconsistent retrieval performance for different image queries. Finally, we combine both retrieved and generated GPS candidates in Geo-verification for location prediction. Experiments on two well-established datasets IM2GPS3k and YFCC4k verify the superiority of G3 compared to other state-of-the-art methods. Our code [^1] https://github.com/Applied-Machine-Learning-Lab/G3 and data [^2] https://huggingface.co/datasets/Jia-py/MP16-Pro are available online for reproduction.
@@ 2 Related Work
Image Geolocalization. Image Geolocalization is an important task in computer vision zhu2023difftraj ; zhu2023synmob ; zhu2024controltraj , spatial data mining zhao2017incorporating ; zhao2016exploring ; zhang2023promptst ; han2023mitigating , and GeoAI zhao2017modeling ; zhao2022multi ; zhang2023mlpst ; zhang2023autostl . Previous work in image geolocalization can be divided into three main categories: classification-based methods, retrieval-based methods, and generation-based methods. (1) Classification-based methods seo2018cplanet ; vo2017revisiting ; muller2018geolocation ; weyand2016planet ; pramanick2022world ; clark2023we divide the entire earth into multiple grid cells and assign the center coordinates as predicted values. Models are then trained to classify the input image into the correct cell. However, if the actual location of the image is far from the center of the predicted cell, there can still be significant errors, even if the cell prediction is correct. (2) Retrieval-based methods treat the image geolocalization task as a retrieval problem, typically maintaining a database of images workman2015wide ; liu2019lending ; zhu2021vigor ; yang2021cross ; zhu2022transgeo ; tian2017cross ; shi2020looking ; zhu2023r2former or a gallery of GPS coordinates cepeda2023geoclip . They take the most similar images and GPS coordinates to the query image as the predicted values. However, maintaining a global-level image database or GPS gallery is infeasible. (3) Generation-based methods employ large multi-modality models to generate the predicted coordinates for images ligeoreasoner . Zhou et al. zhou2024img2loc introduced retrieval-augmented generation into the geolocalization task and took the retrieved similar images’ coordinates as references to help generate predictions. However, they can not accurately extract visual semantics to indicate an image’s location because they simply use visual similarity to retrieve references and suffer inaccurate prediction when facing heterogeneous query images. In this work, G3 introduces Geo-alignment to incorporate geographical information into image representations to help retrieve similar images in geography and proposes Geo-diversification and Geo-verification to enhance prediction performance and robustness.
Large Multi-modality Models. Inspired by the success of large models jia2018second ; jia2024mill ; li2023agent4ranking in single domains like computer vision wang2024visionllm and natural language processing zhao2023survey , there has been increasing attention on large multi-modality models. CLIP radford2021learning aligns image and text representations through contrastive learning and achieves remarkable model generalization with simple optimization objectives. The Large Language-and-Vision Assistant (LLAVA) liu2024visual effectively combines CLIP’s visual encoder with the powerful language model Vicuna, enhancing the model’s general understanding of both visual and textual information through two-stage instruction tuning. GPT4V achiam2023gpt is a large multimodal model released by OpenAI in 2023, allowing users to input text and images to obtain answers.
@@ 3.1 Database Construction
where $\mathcal{L}_{a,b}$ denotes the loss function of modality $a$ to modality $b$, $\mathbf{e}$ is the encoded representations, and $t$ is the temperature. G3 needs to align image representations with both textual descriptions and GPS data, so the final optimization objective is shown below: $\mathcal{L}=(\mathcal{L}_{\text{image},\text{text}}+\mathcal{L}_{\text{image},\text{gps}}+\mathcal{L}_{\text{text},\text{image}}+\mathcal{L}_{\text{gps},\text{image}})/2$.
@@ 3.2 Location Prediction
Geo-verification. Given the coordinate candidates set $\{c_{1},c_{2},\cdots,c_{m}\}$, selecting the best guess is essential and challenging. We reinvent the well-trained Image-to-GPS model in Geo-alignment to achieve this target, as shown in Figure 2(D). The similarity between image representations $\mathbf{e}_{\text{gps}}^{\text{image}}$ and GPS representations $\mathbf{e}^{gps}$ are derived by $\text{sim}=\mathbf{e}_{\text{gps}}^{\text{image}}(\mathbf{e}^{gps})^{\text{T}}$, and the coordinate $c_{j}$ with the highest similarity is selected as the final prediction by $j=\text{argmax}(\text{sim}_{j}),\ j\in\{1,2,\cdots,m\}$.
@@ 4 MP16-Pro Dataset
To facilitate subsequent research, we propose the MP16-Pro dataset by adding textual geographical descriptions to each sample from the MediaEval Placing Tasks 2016 (MP-16) dataset larson2017benchmarking . Specifically, we utilize the open-source geocoding tool Nominatim to obtain multi-level geographical textual descriptions for each sample’s GPS location (total 4.72 million locations). There are eight geographical unit levels: neighborhood, city, county, state, region, country, country code, and continent. Some examples are given in Appendix A.1 for reference. Geographical text descriptions provide additional information for geolocalization tasks and enable models to transcend the original paradigm solely supporting image and GPS alignment, facilitating more diverse modeling approaches.
@@ 5 Experiments
Implementation details: We use faiss douze2024faiss to deploy the image database. The vision and text encoders are pretrained ViT-L/14 and a masked self-attention transformer from CLIP radford2021learning . The dimensions for two trainable layers of $f_{\text{text}}$, $f_{\text{gps}}$, $f$ are 768 and 768. The input dimension of GPS encoder is 512, and the dimensions for four hidden layers of $f_{k}$ in Equation 3.1 are 1024, the output dimension is 512. For the Earth radius, we set it as 6378137.0. For RFF, we use three hierarchies with $\sigma_{min}$ as $2^{0}$ and $\sigma_{max}$ as $2^{8}$. GPT4V  [^4] https://openai.com/ is selected as the LMMs in this paper. Its temperature is set to 1.2. The number of RAG prompts $K$ is set to 4, and the number of candidates for each RAG $N$ is set to 5 for IM2GPS3K and 1 for YFCC4K. The number of similar image coordinates taken into account in candidates is 0 for IM2GPS3K and 1 for YFCC4K. G3 is trained using AdamW optimizer with learning rate 3e-5 and weight decay 1e-6. A step linear scheduler is employed with gamma 0.87, and the training epoch is set to 10. Training batch size is set to 256 and temperature $t$ in Equation 2 is initialized as 3.99. All experiments are conducted with Pytorch and one NVIDIA H800 GPU. Please refer to Appendix A.2 for more details on the training environment, training time, and API cost. We also mention the limitations of G3 in Appendix A.4.
Baselines: To evaluate G3 in geolocalization, we follow previous work cepeda2023geoclip ; zhou2024img2loc and select the following baselines for comparison: [L]kNN,$\sigma$=4 vo2017revisiting , PlaNet weyand2016planet , CPlaNet seo2018cplanet , ISNs muller2018geolocation , Translocator pramanick2022world , GeoDecoder clark2023we , GeoCLIP cepeda2023geoclip , Img2Loc zhou2024img2loc , PIGEON haas2024pigeon . The detailed descriptions of baselines are in Appendix A.5. Due to the lack of available implementations for Img2Loc, we reproduce it based on its paper and release it in our repository for future research.
@@ 5.1 Comparison with State-of-the-art Methods
Table 1: Overall experimental results on IM2GPS3K and YFCC4K.
@@
[L]kNN, sigma=4 vo2017revisiting 7.2 19.4 26.9 38.9 55.9 2.3 5.7 11 23.5 42
PlaNet seo2018cplanet 8.5 24.8 34.3 48.4 64.6 5.6 14.3 22.2 36.4 55.8
CPlaNet seo2018cplanet 10.2 26.5 34.6 48.6 64.6 7.9 14.8 21.9 36.4 55.5
ISNs muller2018geolocation 10.5 28 36.6 49.7 66 6.5 16.2 23.8 37.4 55
Translocator pramanick2022world 11.8 31.1 46.7 58.9 80.1 8.4 18.6 27 41.1 60.4
GeoDecoder clark2023we 12.8 33.5 45.9 61 76.1 10.3 24.4 33.9 50 68.7
GeoCLIP cepeda2023geoclip 14.11 34.47 50.65 69.67 83.82 9.59 19.31 32.63 55 74.69
Img2Loc zhou2024img2loc 15.34 39.83 53.59 69.7 82.78 19.78 30.71 41.4 58.11 74.07
PIGEON haas2024pigeon 11.3 36.7 53.8 72.4 85.3 10.4 23.7 40.6 62.2 77.7
Ours 16.65 40.94 55.56 71.24 84.68 23.99 35.89 46.98 64.26 78.15
To verify the effectiveness of G3, we conduct comparative experiments on IM2GPS3K and YFCC4K with other state-of-the-art methods. The results are shown in Table 1. (1) G3 is superior to all the other baselines on almost all metrics. In addition, compared to the second best methods, G3 achieves 8.5%, 2.8%, 3.3% improvements on IM2GPS3K in the 1km, 25km, 200km thresholds and 21.3%, 16.9%, 13.5%, 3.3%, 0.6% improvements on YFCC4K in the 1km, 25km, 200km, 750km, 2500km thresholds. (2) G3, Img2Loc, GeoCLIP, and PIGEON achieve leading results, which can be attributed to the other methods taking the worldwide geolocalization task as a classification problem, introducing inevitable systemic biases. (3) G3 demonstrates significant improvements over GeoCLIP because GeoCLIP is constrained by the settings of the GPS gallery, which can not cover the entire globe. (4) Compared to Img2Loc, G3, through Geo-alignment that aligns images with discrete and continuous geographical features, achieves more precise retrieval of reference coordinates for subsequent RAG processes. Additionally, Geo-diversification and Geo-verification effectively expand the candidate pool and filter out the confident prediction results, further enhancing geolocalization performance. Overall, G3 achieves the best performance on all datasets across almost all metrics, which verifies the effectiveness of G3.
@@ 5.4 Effectiveness of Geo-alignment and Mercator Projection
5.5 Impact of LMMs on G3
Table 4: Impact of LMMs on G3.
Methods Street 1km City 25km Region 200km Country 750km Continent 2500km
Img2Loc (LLaVA) 10.21 29.06 39.51 56.36 71.07
Img2Loc (GPT4V) 15.34 39.83 53.59 69.70 82.78
G3 (LLaVA) 14.31 35.87 49.42 66.93 81.78
G3 (GPT4V) 16.65 40.94 55.56 71.24 84.68
To explore the impact of LMMs on G3, we conduct the experiments of G3 and Img2Loc with LLaVA (LLaVA-Next-LLaMA3-8b) on IM2GPS3K. From Table 4 we can find that: (1) After switching LMMs from GPT4V to LLaVA, the performance of G3 shows some decline across various metrics but remains competitive. (2) Additionally, compared to Img2Loc (LLaVA), G3 (LLaVA) significantly outperforms Img2Loc (LLaVA), demonstrating the effectiveness of the proposed modules. (3) Finally, by comparing the performance of G3 equipped with LLaVA and GPT4V to Img2Loc equipped with LLaVA and GPT4V, we can observe that G3 shows more stable performance across different LMMs.
5.6 Necessity Analysis of Three Representations Alignment in Geo-alignment
To verify the necessity of aligning the three representations in Geo-alignment, we conduct experiments of the following variants on IM2GPS3K:
  • IMG: Directly using pre-trained CLIP vision encoder as the encoder.
  • IMG+GPS: Aligning Image representations with GPS representations in Geo-alignment, the textual descriptions are not used.
  • IMG+GPS+TEXT(G3): Aligning three modalities simultaneously in Geo-alignment.
Table 5 shows that: (1) By comparing IMG+GPS+TEXT, IMG+GPS, and IMG, we find that adding GPS and text information can both enhance the feature representation compared to using the original image information alone. (2) By comparing IMG+GPS+TEXT with IMG+GPS, we find that IMG+GPS performs better at smaller scales, while IMG+GPS+TEXT performs better at larger scales. This might be because GPS is suitable for modeling variations at smaller scales, whereas text descriptions do not vary significantly at small scales and may even remain the same.
Table 5: Results of necessity analysis.
Methods Street 1km City 25km Region 200km Country 750km Continent 2500km
IMG 15.71 40.64 54.85 70.8 84.05
IMG+GPS 16.91 41.41 55.02 70.94 84.18
IMG+GPS+TEXT(G3) 16.65 40.94 55.56 71.24 84.68
@@ 6 Conclusion
In this paper, we propose a novel worldwide geolocalization framework named G3. First, we introduce Geo-alignment to capture location-aware semantics in images by aligning images with textual geographical descriptions and GPS information. Second, Geo-diversification is proposed to improve the robustness of prediction generation via a prompt ensemble technique. Finally, Geo-verification selects the final coordinate prediction using the learned multi-modality representations. G3 is evaluated on two well-established datasets, IM2GPS3K and YFCC4K, and achieves state-of-the-art performance. In addition, we release a new dataset MP16-Pro, adding textual localization descriptions to each sample based on the original dataset MP16 to facilitate future research in the field. All the code and data used in this work have been released public.
Acknowledgements
This research was partially supported by Research Impact Fund (No.R1015-23), APRC - CityU New Research Initiatives (No.9610565, Start-up Grant for New Faculty of CityU), CityU - HKIDS Early Career Research Grant (No.9360163), Hong Kong ITC Innovation and Technology Fund Midstream Research Programme for Universities Project (No.ITS/034/22MS), Hong Kong Environmental and Conservation Fund (No. 88/2022), and SIRG - CityU Strategic Interdisciplinary Research Grant (No.7020046), Huawei (Huawei Innovation Research Program), Tencent (CCF-Tencent Open Fund, Tencent Rhino-Bird Focused Research Program), Ant Group (CCF-Ant Research Fund, Ant Group Research Fund), Alibaba (CCF-Alimama Tech Kangaroo Fund (No. 2024002)), CCF-BaiChuan-Ebtech Foundation Model Fund, and Kuaishou.
@@ A.1 MP16-Pro Dataset Samples
All the data has been released online [^5] https://huggingface.co/datasets/Jia-py/MP16-Pro.